OpenStudy (anonymous):

(TCO 6) Suppose that an object is traveling along a straight line. Suppose that its velocity function is given by v(t) = 8t^2 -6t where t is in seconds. Find the object's instantaneous acceleration at t=7 seconds. 93 100 106 118 120

OpenStudy (anonymous):

Compute the derivative of the velocity function with respect to time:\[a(t)=v'(t)=16t-6\]Solve when t=7\[a(7)=16(7)-6=106\]: